One number is 3 more than 2 times another. Their product is 14
Answer by stanbon(75887)

One number is 3 more than 2 times another. Their product is 14
------
y = 2x + 3
xy = 14
------
x(2x+3) = 14
2x^2 + 3x - 14 = 0
---
Factors:
2x^2 + 7x - 4x - 14 = 0
------
2x^2 - 4x + 7x -14 = 0
2x(x-2)+7(x-2) = 0
(x-2)(2x+7) = 0
x = 2 or x = -7/2
--------
If x = 2, y = 7
If x = -7/2 , y = -4
Cheers,
Stan H.